OpenStudy (jenniferjuice):

identify the slope –3x + 2y = 8

OpenStudy (jenniferjuice):

?

OpenStudy (jenniferjuice):

ok thanks!

OpenStudy (amistre64):

its not 3

OpenStudy (amistre64):

the slope of: Ax + By = C is not -A, in general

OpenStudy (amistre64):

ax + by = c , solving for y by = -ax + c y = -a/b x + c/b , equating the m part is -a/b

OpenStudy (jenniferjuice):

...

OpenStudy (jenniferjuice):

so i woud divide -2y/-2 = -3x + 8

OpenStudy (jenniferjuice):

then divide 8/-2

OpenStudy (jenniferjuice):

and then i would have y= -3x - 4

OpenStudy (jenniferjuice):

?

OpenStudy (amistre64):

–3x + 2y = 8 2y = 3x + 8 , that +8 is just extra baggage at the moment, so ignore it for slope stuff 2y = 3x /2 /2 y = 3/2 x

OpenStudy (texaschic101):

do you understand how she got that answer ?

OpenStudy (texaschic101):

you want to put it in y = mx + b where m is the slope -3x + 2y = 8 (you want to add 3x to both sides to get 2y by itself) -3x + 3x + 2y = 3x + 8 (combine like terms) 2y = 3x + 8 (now you want to divide both sides by 2 to get y by itself) 2y/2 = 3/2x + 8/2 y = 3/2x + 4 slope being 3/2